# Exponential separations between classical and quantum learners

### v1

@@ Appendix B Proof of Theorem 3
To see why $L_{\mathrm{modexp}}$ is not in $\mathsf{CC}$, we first note that the modular exponentiation concept class contains the cube root function $f_{N}^{-1}$ discussed in Section 3.2. Therefore, the proof presented in [KV94b], which shows that the cube root concept class is not in $\mathsf{CC}$ under the DCRA, also implies that the modular exponentiation concept class is not in $\mathsf{CC}$ under the DCRA. To briefly recap, recall from Section 3.2 that we can efficiently generate examples $(y,f_{N}^{-1}(y))$, for $y\in\mathbb{Z}_{N}^{*}$ uniformly at random. If we put these examples into an efficient classical learning algorithm for the modular exponentiation concept class, the learning algorithm would with high probability identify a classically efficiently evaluatable hypothesis that agrees with $f_{N}^{-1}$ on a polynomial fraction of inputs. This directly violates the DCRA, which states that evaluating $f_{N}^{-1}$ is classically intractable, even on a polynomial fraction of inputs (i.e., it is outside of $\mathsf{HeurBPP}$).
@@
What remains to be shown is that $L_{\mathrm{modexp}}$ is in $\mathsf{QC}$. Suppose we are given access to an example oracle $EX(c_{d},\mathcal{D}^{U}_{n})$ which when queried returns an example $(x,x^{d}\mod N)$, where $x$ is sampled uniformly at random from $\mathbb{Z}_{N}^{*}$. To show that $L_{\mathrm{modexp}}$ is in $\mathsf{QC}$, we will describe a $\mathcal{O}(\mathrm{poly}(n,1/\delta))$-time quantum learning algorithm that uses queries to $EX(c_{d},\mathcal{D}^{U}_{n})$ and identifies $d$ with probability at least $1-\delta$. Before describing the quantum learning algorithm, we will first prove the following lemma, which is used to prove that our quantum learning algorithm is correct.
@@
$*\displaystyle\frac{2^{c_{1}}\cdot 2^{c_{2}}\cdot p^{k_{1}}_{1}\cdots p^{k_{i-1}}_{i-1}\cdot\varphi(p^{k_{i}}_{i})\cdot p^{k_{i+1}}_{i+1}\cdots p^{k_{\ell}}_{\ell}.}{\#\mathbb{Z}_{N}^{*}}$ $=\frac{2^{c_{1}}\cdot 2^{c_{2}}\cdot p^{k_{1}}_{1}\cdots p^{k_{i-1}}_{i-1}\cdot\varphi(p^{k_{i}}_{i})\cdot p^{k_{i+1}}_{i+1}\cdots p^{k_{\ell}}_{\ell}.}{(p-1)(q-1)}$
@@
B.1 Discrete cube root assumption for moduli of Definition 16
@@
Appendix C Proof of Theorem 4
See 4
@@ Appendix C Proof of Theorem 4
To show quantum learnability, we note that $N$ is known and we can thus use Shor’s algorithm [Sho99] to efficiently compute $d\in\{0,\dots,(p-1)(q-1)\}$ such that
@@
Next, we note that from $(x,c_{m}(x))$ we can retrieve the $k$th bit of $m^{3}$, where $k=\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor})$. Since for any given $k\in[n]$ we have
@@
we find that $\mathcal{O}(\mathrm{poly}(n))$ examples suffices to reconstruct the full binary representation of $m^{3}$ with high probability. Finally, using $d$ and $m^{3}$ we can compute $(m^{3})^{d}\equiv m\mod N$.
@@
$\mathsf{Pr}_{k\sim[n]}\left(\mathrm{bin}(m^{3},k)\neq\mathrm{bin}((m^{\prime})^{3},k)\right)\leq\frac{1}{n^{3}},$ (30)
@@
Appendix D Proof of Theorem 5
See 5
@@ D.2 Proof of Lemma 4
Appendix E Proof of Theorem 6
See 6
@@ Appendix E Proof of Theorem 6
Note that the Hamiltonian family constructed above indeed does not satisfy all requirements for the methods of Huang et al. [${}^{+}$22b] to work. In particular, it is known that the spectral gap of Hamiltonians obtained from Kitaev’s circuit-to-Hamiltonian construction (i.e., those defined in Eq. (41)) have a spectral gap that is inverse polynomial in the depth of the circuit, which is our case is polynomial in the instance size $n$. Moreover, since we apply a function $g_{\gamma}$ to the parameters $x$ (which has a rapid increase between $-\gamma$ and $\gamma$), it is likely that the average gradient of the function $\mathrm{Tr}\left[O\rho_{H}(x)\right]$ is not bounded by a constant, but rather scales with the number of parameters $m$ (which in our case also scales with the instance size $n$).

### v2

@@ Appendix D Proof of Theorem 7
To see why $L_{\mathrm{modexp}}$ is not in $\mathsf{CC}$, we first note that the modular exponentiation concept class contains the cube root function $f_{N}^{-1}$ discussed in Section 3.2. Therefore, the proof presented in [KV94b], which shows that the cube root concept class is not in $\mathsf{CC}$ under the DCRA, also implies that the modular exponentiation concept class is not in $\mathsf{CC}$ under the DCRA. To briefly recap, recall from Section 3.2 that we can efficiently generate examples $(y,f_{N}^{-1}(y))$, for $y\in\mathbb{Z}_{N}^{*}$ uniformly at random. If we put these examples (together with examples for $b=1$) into an efficient classical learning algorithm for the modular exponentiation concept class, the learning algorithm would with high probability identify a classically efficiently evaluatable hypothesis that agrees with $f_{N}^{-1}$ on a polynomial fraction of inputs. This directly violates the $2^{c}$-DCRA, which states that evaluating $f_{N}^{-1}$ is classically intractable, even on an inverse polynomial fraction of inputs.
@@
What remains to be shown is that $L_{\mathrm{modexp}}$ is in $\mathsf{QC}$. First, we note that a quantum learner can with high probability recover the modulus $N$ from $\mathcal{O}\left(1\right)$ examples drawn from $E(c_{(N,d)},\mathcal{D}_{n}^{U})$ (i.e., it just needs one for which $b=1$). Next, we note that $EX(c_{d},\mathcal{D}^{U}_{n})$ will in $\mathcal{O}(1)$ queries with high probability return an example $(x,x^{d}\mod N)$, where $x$ is sampled uniformly at random from $\mathbb{Z}_{N}^{*}$. To show that $L_{\mathrm{modexp}}$ is in $\mathsf{QC}$, we will describe a $\mathcal{O}(\mathrm{poly}(n,1/\delta))$-time quantum learning algorithm that uses queries to $EX(c_{d},\mathcal{D}^{U}_{n})$ and identifies $d$ with probability at least $1-\delta$. Before describing the quantum learning algorithm, we will first prove the following lemma, which we will later use to help us prove that our learning algorithm is both correct and efficient.
@@
$\frac{2^{c_{1}}\cdot 2^{c_{2}}\cdot p^{k_{1}}_{1}\cdots p^{k_{i-1}}_{i-1}\cdot\varphi(p^{k_{i}}_{i})\cdot p^{k_{i+1}}_{i+1}\cdots p^{k_{\ell}}_{\ell}.}{\#\mathbb{Z}_{N}^{*}}$ $=\frac{2^{c_{1}}\cdot 2^{c_{2}}\cdot p^{k_{1}}_{1}\cdots p^{k_{i-1}}_{i-1}\cdot\varphi(p^{k_{i}}_{i})\cdot p^{k_{i+1}}_{i+1}\cdots p^{k_{\ell}}_{\ell}.}{(p-1)(q-1)}$
@@
D.1 Discrete cube root assumption for moduli of Definition 20
@@
Appendix E Proof of Theorem 8
See 8
@@ Appendix E Proof of Theorem 8
For quantum learnability, we first note that a quantum learner can with high probability recover the modulus $N$ from $\mathcal{O}\left(\mathrm{poly}(n)\right)$ examples drawn from the example oracle $E(c_{(N,m)},\mathcal{D}^{U}_{n})$. Next, we note that if $N$ is known, then we can use Shor’s algorithm [Sho99] to efficiently compute the integer $d\in\{0,\dots,(p-1)(q-1)\}$ such that
@@
Next, we note that from examples of the form $(x,\mathrm{bin}(m,k))$ we can retrieve the $k$th bit of $m^{3}$, where $k=\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor})$. Since for any given $k\in[n]$ we have
@@
we find that $\mathcal{O}(\mathrm{poly}(n))$ examples of the form $(x,\mathrm{bin}(m,k))$ suffices to reconstruct the full binary representation of $m^{3}$ with high probability. Moreover, we can with high probability obtain an example of the form $(x,\mathrm{bin}(m,k))$ using just $\mathcal{O}(1)$ queries to the example oracle. Finally, using $d$ and $m^{3}$ we can compute $(m^{3})^{d}\equiv m\mod N$.
@@
$\mathsf{Pr}_{k\sim[n]}\left(\mathrm{bin}(m^{3},k)\neq\mathrm{bin}((m^{\prime})^{3},k)\right)\leq\frac{1}{2n^{3}},$ (39)
@@
Appendix F Proof of Theorem 9
See 9
@@ F.2 Proof of Lemma 4
F.3 Proof of Corollary 5
See 5
Proof.
Suppose there exists $L\in\mathsf{BQP}$-$\mathsf{complete}$ such that for every efficiently samplable distribution $\mathcal{D}$ we have $(L,\mathcal{D})\in\mathsf{HeurBPP}/\mathsf{samp}$. Now let $f$ be the (many-to-one) reduction from $\mathsf{DLP}$ to $L$ and define $\mathcal{D}_{L}=f(\mathcal{D}_{U})$ to be the push-forward distribution of the uniform distribution $\mathcal{D}_{U}$ under $f$. By our assumption at the beginning of the proof, we know $(L,\mathcal{D}_{L})\in\mathsf{HeurBPP}/\mathsf{samp}$. However, we now note that the $\mathsf{HeurBPP}/\mathsf{samp}$-algorithm for solving $(L,\mathcal{D}_{L})$ actually also solves $(\mathsf{DLP},\mathcal{D}_{U})$, which implies that $(\mathsf{DLP},\mathcal{D}_{U})\in\mathsf{HeurBPP}/\mathsf{samp}$. Since $\mathsf{DLP}$ is both random self-reducible as well as random verifiable, we conclude that this implies that $\mathsf{DLP}\in\mathsf{BPP}$. ∎
Appendix G Proof of Theorem 10
See 10
@@ Appendix G Proof of Theorem 10
Note that the Hamiltonian family constructed above indeed does not satisfy all requirements for the methods of Huang et al. [HKT+22b] to work. In particular, it is known that the spectral gap of Hamiltonians obtained from Kitaev’s circuit-to-Hamiltonian construction (i.e., those defined in Eq. (50)) have a spectral gap that is inverse polynomial in the depth of the circuit, which is our case is polynomial in the instance size $n$. Moreover, since we apply a function $g_{\gamma}$ to the parameters $x$ (which has a rapid increase between $-\gamma$ and $\gamma$), it is likely that the average gradient of the function $\mathrm{Tr}\left[O\rho_{H}(x)\right]$ is not bounded by a constant, but rather scales with the number of parameters $m$ (which in our case also scales with the instance size $n$).